OpenStudy (carolina4567):

You are going to design an advertisement for a new polynomial identity that you are going to invent. Your goal for this activity is to demonstrate the proof of your polynomial identity through an algebraic proof and a numerical proof in an engaging way! Make it so the whole world wants to purchase your polynomial identity and can't imagine living without it! You may do this by making a flier, a newspaper or magazine advertisement, making an infomercial video or audio recording, or designing a visual presentation for investors through a flowchart or PowerPoint. You must:

OpenStudy (carolina4567):

Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships WARNING! No identities used in the lesson may be submitted. Create your own. See what happens when different binomials or trinomials are combined. Below is a list of some sample factors you may use to help develop your own identity. (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d) My Work : (ax + b)(cx + d) We will use this to create a polynomial identity acx^2 + adx + bcx +bd acx^2 + (ad+bc)x + bd (ax+b)(cx+d) = acx^2 + (ad+bc)x + bd

OpenStudy (carolina4567):

Now all I need to do is prove the numerical expression

OpenStudy (carolina4567):

Help Me @iGreen @amistre64 @satellite73

OpenStudy (amistre64):

youve already proved it ....

OpenStudy (carolina4567):

I have ?

OpenStudy (amistre64):

youve worked from your factors into its identity .. that constitutes a direct proof.

OpenStudy (amistre64):

if we had stated that (ax+b)(cx+d) = ac x^2 + (ad + bc)x + bd then we would have to show why such a statement was true

OpenStudy (igreen):

@amistre64 It says 'numerical proof'..all I see is variables.. :l

OpenStudy (igreen):

I mean: She said she had to proof the 'numerical expression'.

OpenStudy (amistre64):

it mentions nothing of a numerical proof 'Demonstrate that your polynomial identity works on numerical relationships'

OpenStudy (amistre64):

a demonstration is not a proof

OpenStudy (carolina4567):

Actully it says Your goal for this activity is to demonstrate the proof of your polynomial identity through an algebraic proof and a numerical proof in an engaging way!

OpenStudy (amistre64):

pfft, to many words in the middle to actually reads lol

OpenStudy (carolina4567):

No worrys

OpenStudy (amistre64):

insted of abcd, pick your favorite numbers to play with

OpenStudy (igreen):

Yeah, just input numbers in there. Then distribute..it should be easy.

OpenStudy (amistre64):

and for the record, as i already stated ... there are no 'numerical' proofs. only demonstrations that satisfy proof for those specific values

OpenStudy (carolina4567):

Ok how about 2 and 6

OpenStudy (amistre64):

2 numbers are not 4 numbers, technically speaking

OpenStudy (amistre64):

-1 0 1 2

OpenStudy (carolina4567):

Ok now ?

OpenStudy (amistre64):

now use them .....

OpenStudy (igreen):

He gave you the 4 numbers, now plug them in here: (ax + b)(cx + d)

OpenStudy (carolina4567):

Oh how

OpenStudy (carolina4567):

How would i plug it in @iGreen

OpenStudy (igreen):

A = -1 B = 0 C = 1 D = 2

OpenStudy (carolina4567):

So then ( -1x + 0 ) ( 1x + 2 ) Now ?

OpenStudy (igreen):

It would just be (-x + 0)(x + 2) -x + 0 = ?

OpenStudy (carolina4567):

Would it be 0

OpenStudy (igreen):

No any number + 0 is the same thing it was before. 2 + 0 = 2 123 + 0 = 123 651 + 0 = 651

OpenStudy (igreen):

So -x + 0 = ?

OpenStudy (carolina4567):

-x sorry

OpenStudy (igreen):

Yep. So now we have (-x)(x + 2) Now distribute. What's -x * x? And -x * 2?

OpenStudy (carolina4567):

X

OpenStudy (igreen):

-x * x = ?

OpenStudy (carolina4567):

x

OpenStudy (igreen):

No..you will represent it as an exponent. \(-x \cdot x = -x^2\)

OpenStudy (igreen):

Okay, so what's -2 * x?

OpenStudy (igreen):

2 * -x*

OpenStudy (carolina4567):

2x

OpenStudy (igreen):

Not quite, it's -2x. Remember, a negative times a positive is always a negative. So our final result is \(-x^2 - 2x\).

OpenStudy (carolina4567):

THANK U SO MUCH FOR HELPING ME @iGreen

OpenStudy (igreen):

No problem.